Calculate the pH of a 0.22 M solution of hydrocyanic acid (HCN), which has a pKa of 9.31.
1
Expert's answer
2019-03-27T04:35:35-0400

V = 1L.

 

Xmol  Xmol   Xmol

 HCN ⇄ H+ +CN-

 1mol  1mol 1mol

pKa = 9.31;   Ka = 10-9.31 = 4.90×10-10.

Ka = [H+][CN-]/[HCN] = X2/(0.22-X) = 4.90×10-10.

X2 = 4.90×10-10(0.22-X)

X2+4.9×10-10×X–1.08×10-10=0

D = b2 – 4ac = (4.9×10-10)2 – 4×1×(–1.08×10-10) = 4.32×10-10.

"X_1=(-b-\u221aD)\/2a";

X1 = (–4.9×10-10– √4.32×10-10)/2 = –1.04×10-5.

"X_2=(-b+\u221aD)\/2a";

X2 = (–4.9×10-10+√4.32×10-10)/2 = 1.04×10-5.

X1 does not match the condition of the question.

[H+] = X2 = 1.04×10-5.

pH = -lg[H+] = -lg(1.04×10-5) = 4.9833 ≈ 4.98.